OpenStudy (mackenzie2013):

A spinner is divided into 4 equal sections numbered 1 to 4. The theoretical probability of the spinner stopping on 3 is 25%. Which of the following is most likely the number of 3's spun in 10,000 spins? 3,108 2,538 2,400 1,367

OpenStudy (anonymous):

\[10,000*0.25=2500\]\[3108-2500=608\]\[2538-2500=38\]\[2500-2400=100\]\[2500-1367=1133\] 2538 is the most likely.